# Staircase Cascaded Fusion of Lightweight Local Pattern Recognition and Long-Range Dependencies for Structural Crack Segmentation

### v2

@@ Abstract
Detecting cracks with pixel-level precision for key structures is a significant challenge, as existing methods struggle to effectively integrate local textures and pixel dependencies of cracks. Furthermore, these methods often possess numerous parameters and substantial computational requirements, complicating deployment on edge control devices. In this paper, we propose a staircase cascaded fusion crack segmentation network (CrackSCF) that generates high-quality crack segmentation maps using minimal computational resources. We constructed a staircase cascaded fusion module that effectively captures local patterns of cracks and long-range dependencies of pixels, and it can suppress background noise well. To reduce the computational resources required by the model, we introduced a lightweight convolution block, which replaces all convolution operations in the network, significantly reducing the required computation and parameters without affecting the network’s performance. To evaluate our method, we created a challenging benchmark dataset called TUT and conducted experiments on this dataset and five other public datasets. The experimental results indicate that our method offers significant advantages over existing methods, especially in handling background noise interference and detailed crack segmentation. The F1 and mIoU scores on the TUT dataset are 0.8382 and 0.8473, respectively, achieving state-of-the-art (SOTA) performance while requiring the least computational resources. The code and dataset is available at https://github.com/Karl1109/CrackSCF.
@@
Structural cracks segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
@@ I Introduction
Cracks are common defects in structures such as pavements and buildings, which can lead to serious safety issues. Thus, automated detection on edge control devices is crucial for ensuring production safety [1, 2]. Early automated detection methods used traditional image processing techniques for crack extraction [3, 4, 5]. While these methods are easy to implement, their performance in complex scenarios still needs improvement.
@@
The success of VIT [14] and Swin Transformer [15] has demonstrated the advantages of Transformer [16], with their strong sequence processing capabilities, in handling visual tasks. Unlike CNNs, transformers can capture dependencies between any pixels when processing images [17, 18], making them more advantageous in handling complex topologies. To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as CATransUNet [19], MFAFNet [20], and Crackmer [9], which integrate local texture information and long-range pixel dependencies through U-shaped connections or residual connections.
@@
To address these issues, we propose a staircase cascaded fusion network that can generate high-quality segmentation maps with minimal computational resources while effectively suppressing background noise. Specifically, to better exploit the complementarity and correlation of feature maps without missing critical detail information, we introduce a staircase cascaded fusion module. This module utilizes the pixel attention mechanism, which finely captures local and long-range information, and the channel concatenation mechanism, which progressively integrates features from adjacent layers, to generate crack segmentation maps that comprehensively combine local texture information and pixel dependencies. Furthermore, to decrease the parameter count and computational burden, we introduce a lightweight convolution block called LRDS. This block initially increases and then reduces the channel count, and separates the spatial and channel aspects, thereby minimizing computational costs while maintaining performance.
To assess the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images. Existing public datasets like DeepCrack [21] and Crack500 [22] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance. The TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines. With such a diverse set of image scenarios, this dataset effectively assesses the network’s generalization and robustness. Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
@@
    We proposed a novel CrackSCF network for structural crack segmentation. By introducing a staircase cascaded fusion module and a lightweight convolution block, this network effectively distinguishes between background and crack pixels with minimal computational resources, resulting in high-quality segmentation maps.
@@ II-A Crack Segmentation Networks
Although the aforementioned methods have demonstrated the advantages of CNNs in crack segmentation tasks, CNNs are essentially local operations and cannot effectively perceive a small number of crack pixels amidst a large background. Therefore, these methods often result in discontinuous crack pixels and are highly susceptible to background noise interference.
Figure 2: General Architecture Diagram of CrackSCF Network. Crack images are first input into the MFE, generating four feature maps. After enhancement, five feature maps with unified channel numbers are obtained. The LDE module processes these maps to acquire pixel sequences rich in long-range correlations. The SCFM module then processes these sequences and the four layers of feature maps from the MFE. Through four stages, resulting in the segmentation output.
Thanks to the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes. Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation. Chen et al. [25] proposed TransUNet, a novel network architecture that effectively integrates long-range information and local features through a U-Net-like structure. Qi et al. [26] developed an end-to-end model based on Vision Transformer (ViT) and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders. Feng et al. [27] developed a shale crack segmentation network that combines CNN and Swin Transformer, utilizing Fast Fourier Transform (FFT) to merge feature maps from both, effectively suppressing high-frequency noise in the imaging process.
@@ II-B Lightweight Networks
Deploying deep learning models on edge devices like smartphones and drones is trending. However, these devices have limited computing resources, making it challenging to use models with high parameter counts. Thus, compact model design for lightweight modifications in size and speed is crucial. For example, Howard et al. [28] proposed MobileNet, an efficient model using depthwise separable convolutions to reduce computational load and parameter count, suitable for resource-constrained devices. Ma et al. [29] introduced ShuffleNet, providing practical design principles for efficient network architecture. Additionally, low-rank approximation compression, effective in fields like numerical linear algebra and image processing, is now used to compress deep learning network architectures. Zhang et al. [30] proposed a low-rank technique for neural networks that employs constrained optimization to to determine the optimal low-rank approximation. Swaminathan et al. [31] introduced SLR, a sparse low-rank method that improves compression rates with sparse SVD matrices.
Recently, many researchers have combined model lightweighting techniques with crack segmentation models to enable real-time crack detection on edge devices. For example, Zhu et al. [32] designed RHACrackNet, a lightweight crack detection network integrating hybrid attention and residual blocks, enabling real-time detection on edge devices. Huang et al. [33] proposed a lightweight fusion attention module for road crack detection, ensuring high-precision crack extraction with rapid detection.
Although these lightweight models reduce computational resources, their emphasis on minimizing parameters results in limited model expressiveness, hindering their ability to effectively capture intricate patterns and features in the input data. This leads to unsatisfactory segmentation results. This imbalance between performance and lightweight design makes them unreliable in practical applications.
@@ III-A Network Architecture
As shown in Fig. 2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM). The overall architecture of the network is shaped like a staircase, with feature maps of different scales progressively fused and upsampled. In this process, the advantages of CNNs and Transformers in feature extraction complement each other.
@@ III-B Lightweight Multi-scale Feature Extractor
As shown in Fig. 2 and Fig. 4, our designed LRDS block adopts a bottleneck-like structure. It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer. The input feature map first passes through the Reduce layer, reducing the input channels to $C_{m}$. It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged. Finally, the Expand layer increases the number of channels to $C_{d}$, producing the final output feature map.
@@
Besides low-rank approximation, which reduces convolutional layer parameters and computational load, depthwise separable convolutions [28] also significantly reduce computational resources by decoupling spatial and channel dimensions. Therefore, we embedded depthwise and pointwise convolution layers in the LRDS block. Additionally, to better preserve local contextual information during feature extraction while avoiding excessive computational overhead, we incorporated strip pooling [34] technology into each bottleneck of ResNet50.
Figure 4: Illustration of LRDS convolution block.
We constructed MFE by replacing all convolution operations in ResNet50 with LRDS convolution blocks.
III-C Long-range Dependency Extractor
@@
When extracting long-range dependencies of crack pixels, the deformable attention mechanism [35] adapts well to irregular structures. It calculates attention weights within a few sampled points, significantly reducing computational and memory overhead. Specifically, given an input $x\in\mathbb{R}^{C\times H\times W}$, let $q$ denote a query element with query feature $z_{q}$ and reference point coordinates $(p_{qx},p_{qy})$, the deformable attention is calculated as:
@@
The structure of LDE is shown in Fig. 2. Specifically, before the enhanced feature maps enters the LDE encoder, it is flattened into a long pixel sequence, aiding the deformable attention mechanism in extracting pixel dependencies. As shown in Fig. 5(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module. To reduce these, we designed a lightweight Linear layer (LR Linear) using low-rank approximation, as shown in Fig. 2. We split the original Linear operation into two parts: $f_{m}$ passes through the Reduce layer to get $f_{mid}$ with 4 channels, then through the Expand layer to produce $f_{out}$. Replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M, as shown in Fig. 5(b).
@@ III-D Staircase Cascaded Fusion Module
Figure 6: Illustration of the PAF and SeFu.
The design of our SCFM structure is shown in Fig. 2. To enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, we improved the Pixel Attention Guided Fusion (PAF) module, inspired by PIDNet [36] and the need for model lightweighting. The structure of PAF is shown in Fig. 6(a). We denote the pixels in the feature maps of the two branches as $\vec{v_{c}}$ and $\vec{v_{d}}$, respectively. Thus, the output can be expressed as:
@@
where $f_{c}$ and $f_{d}$ denote pointwise convolution and batch normalization operations, respectively. $\sigma$ indicates the probability that the two pixels belong to the same category. If $\sigma$ is high, the model will trust $\vec{v_{c}}$ more, indicating that its semantic features are more accurate, and vice versa. Thus, the output of PAF can be represented as::
@@
To improve inference speed and reduce computational resource usage, all convolution operations in this module utilize LRDS blocks. As seen in Fig. 3, using original convolution operations in the SCFM results in FLOPs and Params reaching 73.36G and 9.71M, respectively. Thus, lightweighting these operations is crucial. Replacing all convolution operations with LRDS blocks reduces FLOPs and Params by 92.87% and 97.22%, respectively, demonstrating the effectiveness of LRDS in network lightweighting. SCFM is divided into four stages, each with two fusion branches: the concat branch and the PAF branch. The concat operation retains complete information and diverse feature representations through direct concatenation, while the PAF module selectively fuses features through a pixel-guided attention mechanism, avoiding issues like buried feature information, information conflict, and redundancy. By perceiving the feature maps from the two branches, both local patterns of crack pixels and long-range dependencies can be comprehensively addressed.
To fully utilize the outputs of MFE and LDE and improve segmentation performance, SCFM incorporates the outputs from earlier stages and the associated MFE feature maps starting from the second stage. These are fused through the PAF module and processed by the Selective Feature Fusion (SeFu) module [7], which combines all PAF outputs before further operations. This approach effectively combines local semantic and long-range information, improving segmentation. The SeFu module structure is shown in Fig. 6(b).

### v3

@@ Abstract
Detecting cracks with pixel-level precision for key structures is a significant challenge, existing methods struggle to integrate local textures and pixel dependencies of cracks. Furthermore, these methods possess numerous parameters and substantial computational requirements, complicating deployment on edge devices. In this paper, we propose the Staircase Cascaded Fusion Crack Segmentation Network (CrackSCF), which generates high-quality crack segmentation maps while reducing computational overhead. We design a lightweight convolutional block that substitutes all convolution operations, reducing the model’s computational demands while maintaining an effective capture of local details. Additionally, we introduce a lightweight long-range dependency extractor to better capture the long-range dependencies. Furthermore, we develop a staircase cascaded fusion module, which seamlessly integrates local patterns and long-range dependencies, resulting in high-quality segmentation maps. To comprehensively evaluate our method, we created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets. The results show that our method outperforms existing methods, particularly in handling background noise and achieving detailed segmentation. The F1 and mIoU scores on the TUT dataset are 0.8382 and 0.8473, respectively, demonstrating state-of-the-art (SOTA) performance with low computational resources. The code and dataset is available at https://github.com/Karl1109/CrackSCF.
@@
Structural crack segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
publicationid: pubid: $0$–$00$ 2025 IEEE
@@ I Introduction
Cracks are common defects in structures such as pavements and buildings, which can lead to serious safety issues. Thus, automated detection on edge control devices is crucial for ensuring production safety [1]. Early automated detection methods used traditional image processing techniques for crack extraction [2, 3, 4]. While these methods are easy to implement, their performance in complex scenarios still needs improvement.
@@
The success of Vision Transformer (ViT) [13] and Swin Transformer [14] has demonstrated the advantages of Transformer [15], with their strong sequence processing capabilities, in handling visual tasks [16, 17]. Unlike CNNs, transformers can capture dependencies between any pixels when processing images [18, 19], making them more advantageous in handling complex topologies. To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as CATransUNet [20], MFAFNet [21], and Crackmer [8], which integrate local and long-range information through U-shaped connections or residual connections.
@@
To address these issues, we propose a staircase cascaded fusion network that generates high-quality segmentation maps with low computational resources while effectively suppressing background noise. Specifically, to reduce the number of parameters and computational burden, we introduce a lightweight convolution block called LRDS. This module utilizes the principle of low-rank approximation to effectively reduce the computational resources required by convolutional layers, and captures spatial structures efficiently in the low-rank space through depthwise convolution, maintaining good feature perception while requiring very low computational resources. We also propose a lightweight Long-range Dependency Extractor (LDE), which uses a lightweight deformable attention mechanism to effectively capture irregular long-range dependencies between crack pixels with very low computational demands. Additionally, to better leverage the complementarity and correlation of feature maps while ensuring that critical detail information is not missed, we introduce the Staircase Cascaded Fusion Module (SCFM). This module generates feature maps layer by layer and effectively fuses them using a pixel attention mechanism that can finely capture local and long-range information and a channel concatenation mechanism, comprehensively combining local texture information and pixel dependencies.
To comprehensively evaluate the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images. Existing public datasets like DeepCrack [22] and Crack500 [23] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance. The TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines. With such a diverse set of image scenarios, this dataset effectively evaluates the network’s generalization and robustness. Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
@@
    We proposed a novel CrackSCF network for structural crack segmentation. By introducing a lightweight convolution block, a lightweight long-range dependency extractor, and a staircase cascaded fusion module, this network effectively distinguishes between background and crack pixels with low computational resources, resulting in high-quality segmentation maps.
@@ II Related Work
Figure 2: General Architecture Diagram of CrackSCF Network. Crack images are first input into the MFE, generating four feature maps. After enhancement, five feature maps with unified channel numbers are obtained. The LDE processes these maps to acquire pixel sequences rich in long-range correlations. The SCFM then processes these sequences and the four layers of feature maps from the MFE. Through four stages, resulting in the segmentation output.
@@ II-A Crack Segmentation Networks
Although the aforementioned methods have demonstrated the advantages of CNNs in crack segmentation tasks, CNNs are essentially local operations and cannot effectively perceive a small number of pixels amidst a large background. Therefore, these methods often result in discontinuous crack pixels and are highly susceptible to background noise interference.
Benefit from the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes. Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation. Chen et al. [26] proposed TransUNet, a novel network that effectively integrates long-range information and local features through a U-Net-like structure. Qi et al. [27] developed an end-to-end model based on ViT and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders. Feng et al. [28] developed a shale crack segmentation network that combines CNN and Swin Transformer, utilizing Fast Fourier Transform (FFT) to merge feature maps from both, effectively suppressing high-frequency noise in the imaging process.
@@ II-B Lightweight Networks
Deploying deep learning models on edge devices like smartphones and drones is trending. However, these devices have limited computing resources, making it challenging to use models with high parameter counts. Thus, compact model design for lightweight modifications in size and speed is crucial [29, 30]. For example, Howard et al. [31] proposed MobileNet, an efficient model using depthwise separable convolutions to reduce computational load and parameter count, suitable for resource-constrained devices. Ma et al. [32] introduced ShuffleNet, providing practical design principles for efficient network architecture. Additionally, low-rank approximation compression, effective in fields like numerical linear algebra and image processing, is now used to compress deep learning network architectures. Swaminathan et al. [33] introduced SLR, a sparse low-rank method that improves compression rates with sparse SVD matrices.
In recent years, there has been a growing trend among researchers to integrate model lightweighting techniques with crack segmentation models, aiming to facilitate real-time crack detection on edge computing devices. For example, Zhu et al. [34] designed RHACrackNet, a lightweight crack detection network integrating hybrid attention and residual blocks, enabling real-time detection on edge devices. Huang et al. [35] proposed a lightweight fusion attention module for road crack detection, ensuring high-precision crack extraction with rapid detection. Although these lightweight models reduce computational resources, their emphasis on minimizing parameters results in limited model expressiveness, hindering their ability to effectively capture intricate patterns and features in the input data. This leads to unsatisfactory segmentation results. This imbalance between performance and lightweight design makes them unreliable in practical applications.
@@ III-A Network Architecture
As shown in Fig. 2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a lightweight Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM). The overall architecture of the network is shaped like a staircase, with feature maps of different scales progressively fused and upsampled. In this process, the advantages of CNNs and Transformers in feature extraction complement each other.
@@ III-B Lightweight Multi-scale Feature Extractor
As shown in Fig. 4, our designed LRDS block adopts a bottleneck-like structure. It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer. The input feature map first passes through the Reduce layer, reducing the input channels to $C_{m}$. It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged. Finally, the Expand layer increases the number of channels to $C_{d}$, producing the final output feature map.
@@
Low-rank approximation theory greatly reduces the convolutional layer parameters and computational load. In order to efficiently capture the spatial information of feature maps in low-rank space for finer perception of detailed regions, we process the features in LRDS by embedding deep convolution and pointwise convolution in low-rank space. We constructed MFE by replacing all convolution operations in ResNet50 with LRDS convolution blocks.
Figure 4: Illustration of LRDS convolution block. After the number of channels of the input features is reduced, their local and spatial features are efficiently extracted by depth-wise convolution and point-wise convolution, and finally the number of channels is restored.
III-C Lightweight Long-range Dependency Extractor
@@
When extracting long-range dependencies of crack pixels, the deformable attention mechanism adapts well to irregular structures. Specifically, given an input $x\in\mathbb{R}^{C\times H\times W}$, let $q$ denote a query element with query feature $z_{q}$ and reference point coordinates $(p_{qx},p_{qy})$, the deformable attention is calculated as:
@@
As shown in Fig. 5(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module. In order to reduce the computational resources required by the LDE, we propose a linear transformation method (LR Linear) based on low-rank approximation. In traditional linear layers, the weight matrix connecting the input and output is typically a high-rank matrix, representing a complete connection between input and output channels. By applying low-rank approximation, we decompose the weight matrix into two smaller matrices, reducing the number of parameters.
Specifically, the weight matrix $E\in\mathbb{R}^{r_{\text{out}}\times r_{\text{in}}}$, where $r_{\text{out}}$ is the output feature dimension and $r_{\text{in}}$ is the input feature dimension, is approximated by two smaller matrices $A\in\mathbb{R}^{r_{\text{out}}\times r}$ and $B\in\mathbb{R}^{r_{\text{in}}\times r}$, where $r$ is a small rank, much smaller than $r_{\text{out}}$ and $r_{\text{in}}$. This decomposition breaks the original full connection into two smaller linear transformations:
$E\approx E^{\prime}=AB^{T}$ (5)
we perform this transformation through two linear layers. First, we apply a linear transformation $B^{T}$ to the input $x\in\mathbb{R}^{r_{\text{in}}}$, resulting in an intermediate representation $z\in\mathbb{R}^{r}$:
$z=B^{T}x$ (6)
Then, we apply another linear transformation $A$ to $z$, resulting in the output $y\in\mathbb{R}^{r_{\text{out}}}$:
$y=Az$ (7)
As $r\ll r_{\text{out}},r_{\text{in}}$, the low-rank approximation reduces the original matrix to two smaller matrices, significantly reducing the computational and storage requirements. This leads to a reduction in both the computational complexity and the number of parameters, thus improving training and inference efficiency. As shown in Fig. 5(b), replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M.
The structure of LDE is shown in Fig. 2. LDE facilitates long-range dependency modeling through the interaction of multi-scale features between the encoder and decoder. Initially, the encoder inputs a multi-scale feature set $\{X_{c}\}_{c=1}^{C}$ (where $C$ represents the number of feature levels). After undergoing layer flattening and positional encoding enhancement, the encoder extracts the feature sequence $S\in\mathbb{R}^{B\times L\times d}$ (where $L$ denotes the sequence length). The encoder then combines LR Linear with lightweight deformable attention mechanisms to capture long-range dependencies:
$Z=\text{MSDeformAttn}(S,P,S)$ (8)
where $P\in\mathbb{R}^{B\times L\times 2}$ represent the normalized reference coordinates. Subsequently, $Z$ is processed through a feedforward network (FFN) that combines LR Linear to refine the features:
$M=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Z)))$ (9)
where the output memory matrix $M\in\mathbb{R}^{B\times L\times d}$ is used as the input to the decoder, encoding multi-scale local dependencies.
The decoder processes the input with a learnable query vector $Q\in\mathbb{R}^{N\times d}$ and the initial reference point $H=\text{Sigmoid}(TQ)$, where $T\in\mathbb{R}^{d\times 2}$ is the projection matrix. First, the reference points $H\in\mathbb{R}^{N\times 2}$ are mapped to the encoded memory $M$ to aggregate features:
$Y_{0}=\text{MSDeformAttn}(Q,H,M)$ (10)
Then, it passes through the FFN to obtain the crack pixel sequence enriched with irregular long-range dependencies:
$Y=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Y_{0})))$ (11)
where Y is the output sequence of the LDE enriched with irregular long-range dependencies between pixels.
@@ III-D Staircase Cascaded Fusion Module
Figure 6: Illustration of the PAF. It effectively reduces required computing resources by embedding LRDS.
The design of our SCFM structure is shown in Fig. 2. As a segmentation head, to enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, we proposed the lightweight Pixel Attention Guided Fusion (PAF) module. The structure of PAF is shown in Fig. 6. We denote the pixels in the feature maps of the two branches as $\vec{v_{c}}$ and $\vec{v_{d}}$, respectively. The output can be expressed as:
@@
where $f_{c}$ and $f_{d}$ denote pointwise convolution and batch normalization operations, respectively. Thus, the output of PAF can be represented as:
@@
where $\sigma$ indicates the probability that the two pixels belong to the same category. If $\sigma$ is high, the model will trust $\vec{v_{c}}$ more, indicating that its semantic features are more accurate, and vice versa.
To reduce computational resource usage, all convolution operations in this module utilize LRDS blocks. As seen in Fig. 3, using original convolution operations in the SCFM results in FLOPs and Params reaching 73.36G and 9.71M, respectively. Thus, lightweighting these operations is crucial. Replacing all convolution operations with LRDS blocks reduces FLOPs and Params by 92.87% and 97.22%, respectively, demonstrating the effectiveness of LRDS in network lightweighting.
SCFM is divided into four stages, each with two fusion branches: the concat branch and the PAF branch. The concat operation retains complete information and diverse feature representations through direct concatenation, while the PAF module selectively fuses features through a pixel-guided attention mechanism, avoiding issues like buried feature information, information conflict, and redundancy. The processes of SCFM can be given by the following equation:
$\begin{split}&F_{i}^{\text{O}}=C_{\text{i}}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\right)\odot P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\end{split}$ (14)
Where $F_{i}^{\text{O}}$ represents the output features of the $i$-th SCFM layer ($i$ ranges from 0 to 3), $F_{i}^{\text{MFE}}$ represents the $i$-th layer MFE features, $C_{\text{i}}$ is an operation chain: Concat followed by LRDS, ReLU, and GN, $P_{i}$ is a PAF operation, $\mu(\cdot)$ is the sigmoid function generating spatial attention weights, and $\odot$ represents the Hadamard product. Specifically, when $i=0$, $F_{i-1}^{\text{O}}$ is defined as the output of LDE, i.e., $F_{-1}^{\text{O}}=\text{LDE}_{\text{output}}$, ensuring an initial output feature is provided at the beginning of the computation.
By perceiving the feature maps from the two branches, both local patterns of crack pixels and long-range dependencies can be comprehensively addressed. To fully utilize the outputs of MFE and LDE and improve segmentation performance, SCFM utilized the property that PAF can process multiple feature maps simultaneously by merging the outputs of all the preceding phases and the associated MFE feature maps from the second phase onwards. The processing enables SCFM to more comprehensively dissect and process the local patterns and long-range dependencies in the different stages of the feature map to achieve a finer and more continuous segmentation map. Furthermore, the fine-grained features extracted with PAF can be perceived more finely through the dynamic weighting mechanism to achieve a clearer segmentation map.